The separation between the consecutive dark fringes in a Young’s double of 2.5 m from the slits and the separation between the slits is 1.0 mm. Calculate the wavelength of light used for the experiment.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. Given that, β = 1 mm = 10^–3 m, D = 2.t m and d = 1 mm = 10^–3 m So, 10^-3m = 25 * λ/10^-3 ⇒ λ = 4 * 10^-7 m = 400 nm.